Friendly’s Quick Loans, Inc., offers you $8.00 today but you must repay $9.95 when you get your paycheck in one week (or else).

What is the effective annual return Friendly’s earns on this lending business? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)

Effective annual return ??????? %
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

9.95=8*(1+r/100)

(9.95/8)=(1+r/100)

(1+r/100)=1.24375

r=1.24375-1

=0.24375

EAR=(1+APR/m)^m-1
where m=compounding periods

=(1+0.24375)^52-1

which is equal to

=8,435,595.12%(Approx).